# Importance-Weighted Non-IID Sampling for Flow Matching Models

### v1

@@ 3.3 Discussions
Since $r_{\phi}$ is trained, one can sample independently from $v+r_{\phi}$ (the ODE in Definition˜3), which matches the marginal distribution of diversity-enhanced samples. However, IID batches drawn from $v+r_{\phi}$ are typically less diverse than those generated by the coupled ODEs in Definition˜2, which explicitly enhance diversity across concurrent trajectories (see Table 1).
@@
In Theorem˜1, we estimate the importance weight at the fixed position for each sample as
@@
which, given $v$ and $v+r_{\phi}$ are rectified flows, further simplifies (Proof in Sec.˜C.2):
@@
However, in practice, the pair $(X_{1}^{(i)},t)$ might be unlikely to occur along the real sampling path when $t<1$, because $X_{1}^{(i)}$ is the sample for time $t=1$. In conditional generation, i.e. generating an image for the text “a cat”, $\mathbb{E}X_{1}^{(i)}\neq\mathbf{0}$ but $\mathbb{E}X_{0}^{(i)}=\mathbf{0}$, which means $X_{1}^{(i)}$ can be rare among real samples at $t=0$, i.e. $p_{0}(X_{1}^{(i)})\approx 0$. This implies that the training data of $r_{\phi}$ cannot cover this situation, potentially leading to an undefined output $r_{\phi}(X_{1}^{(i)},t)$ for small $t$ and worsening performance. Our trajectory-based estimator instead integrates along the actual path and avoids out-of-distribution inputs to $r_{\phi}$ (see Table 2).
@@
Table 1: Sampling diversity and quality results for the Gaussian mixture experiment. Mode coverage, $\log p$, RMSE are reported as mean(uncertainty) with 95% confidence intervals. The best-performing method for each metric is shown in bold.
@@
JS divergence 0.073(1) 0.082(2) 0.512(3) 0.526(3) 0.458(3) 0.087(1) 0.077(1)
Table 3: Jensen-Shannon divergence results for estimation. Values are mean(uncertainty) with 95% confidence intervals. indicates the variation of our method that estimates the importance-weight evolution at a fixed, final sample position. “Equal” means using non-IID samples but with equal weights, while “IID” means using additional IID samples. Lower is better; the best method is in bold.
Method SR $T_{1}$ $T_{2}$ $T_{3}$ $T_{4}$ $T_{5}$ $T_{6}$ Mean
@@
Table 4: Representation error results on text-to-image generation. Values are mean(uncertainty) with 95% confidence intervals.
4 Experiments
We first evaluate on a Gaussian mixture with known density, enabling precise diagnosis of diversity, quality, and estimation of importance weights and expectations. We then assess plug-in benefits on large flow models—Stable Diffusion 3.5 Medium [DBLP:conf/icml/EsserKBEMSLLSBP24] for text-to-image generation and FLUX.1-Fill-dev [DBLP:journals/corr/abs-2506-15742] for image inpainting—to study distribution coverage in complex image settings. Diversity velocities and implementation details for each experiment are in Appendix˜D and Appendix˜E, respectively.
4.1 Experiments on Gaussian Mixture
4.1.1 Sample Diversity and Quality
Setup.
Results.
In Table 1, across all diversity objectives, adding SR (soft or hard) significantly improves quality ($\log p$ increases; RMSE decreases) while preserving joint mode coverage. For example, with DPP, $\log p$ and RMSE improve from $7.16$ and $0.527$ to $19.27$ and $0.253$ under SR-hard, while joint coverage remains almost unchanged ($8.55{\to}8.57$). By contrast, DiverseFlow (DPP+DF) improves quality but reduces diversity ($8.55{\to}6.89$), evidencing a diversity–quality trade-off that our score-based regularization alleviates. Finally, IID sampling from the diversity-enhanced marginal fails to recover joint-diversity benefits (e.g., DPP joint coverage $8.55$ vs. marginal $6.51$, close to original IID $6.51$), justifying our design choice of joint sampling via coupled ODEs (Definition˜2) rather than IID from the diversity-enhanced marginal (Definition˜3).
@@ 4.2 Importance Weight & Expectation Estimation
For accurate ground truth, we use a 2D Gaussian mixture where each component $k$ has a weight proportional to $2^{k}$. Non-IID joint sets are sampled using a Harmonic DPP with soft SR. The ground-truth importance weights are computed as the ratio between the original density and the non-IID marginal estimated via Local-Likelihood Density Estimation (LLDE) [Loader1996, HjortJones1996] (see Appendix˜F). We train the residual velocity $r_{\phi}$ on 1,0001{,}000 joint sets (each containing $10$ trajectories) and evaluate on 10,00010{,}000 sets. We compare three types of estimators: (i) trajectory-based evolution (ours, Theorem˜2); (ii) fixed-position evolution at a terminal state (Theorem˜1); and (iii) density baselines—k-Nearest Neighbors Density Estimation (kNN) [loftsgaarden1965nonparametric], Gaussian Kernel Density Estimation (KDE) [DBLP:books/sp/Silverman86], and Multivariate Gaussian Fit (MGF) [DBLP:journals/jei/BishopN07] (see Appendix˜G).
@@
We report squared error (SE) and ranking metrics—Kendall’s $\tau_{b}$ [kendall1945treatment], Spearman’s $\rho$ [spearman1987proof], and graded AP (gAP) [DBLP:conf/sigir/MoffatM0A22] (see Appendix˜H)–computed on the top $50\%$ most-confident ground truth. In Table 2, density baselines exhibit large SE and weak ranking, while the trajectory-based estimator consistently outperforms the fixed-position variant, since path integration avoids out-of-distribution inputs to $r_{\phi}$.
@@
We estimate $\mu{=}\mathbb{E}f(X)$, where $f(x)$ maps each sample to a one-hot vector indicating the index of its nearest Gaussian. Ground-truth $\mu$ is computed from 1,000,0001{,}000{,}000 IID samples drawn from the original flow-matching model, yielding the probability that a generated sample belongs to each Gaussian. We evaluate on joint samples via Jensen–Shannon (JS) divergence [DBLP:journals/tit/Lin91], which remains well-defined when some predicted probabilities are zero. In Table 3, non-IID samples without weighting are biased and perform worse than IID; fixed-position weighting reduces bias but still trails IID; only the trajectory-based estimator surpasses IID.
@@
Table 5: Representation error results on image inpainting. Values are mean(uncertainty) with 95% confidence intervals.
@@ 4.3 Experiments on Text-to-Image Generation
We next test Stable Diffusion 3.5 Medium [DBLP:conf/icml/EsserKBEMSLLSBP24] for text-to-image generation. We study whether using diverse samples can better cover the IID samples $S_{\text{IID}}$ with a fixed budget of samples $S_{\text{Rep}}$, measured by the representation error metric
@@
Table 4 shows that introducing diversity terms reduces the representation error for all prompts $T_{1}$ to $T_{6}$ relative to IID (taking $S_{\text{Rep}}$ as 10 IID samples), indicating improved coverage under a fixed budget. Adding our score-based regularization (SR) further improves performance across all prompts, suggesting that maintaining on-manifold quality increases sample efficiency (fewer “wasted” samples). Qualitative results are shown in Figure˜3.
Figure 4: Qualitative results for image inpainting. Two samples are shown from the joint samples, with all methods sharing identical initialization. In Sample 1, DPP introduces artifacts (highlighted by the black rectangle) that are removed by SR (hard). While enhancing quality, DPP+SR (hard) retains the diversity of DPP, as illustrated in Sample 2.
@@ 4.4 Experiments on Image Inpainting
We additionally evaluate performance on image inpainting, which is more constrained than text-to-image, to test whether improvements persist. We use FLUX.1-Fill-dev [DBLP:journals/corr/abs-2506-15742]. We select 10 square, high-quality images from the Open Images Dataset V7 [DBLP:journals/ijcv/KuznetsovaRAUKP20] and apply a mask with four squares to inpaint. For each image, we generate $|S_{\text{IID}}|\!=\!1000$ inpaintings and evaluate $|S_{\text{Rep}}|\!=\!10$ jointly sampled inpaintings (100 repetitions). The representation error (Equation˜26) is computed over average latent features within the masked regions.
Table 5 shows that diversity terms reduce representation error for most inputs and improve the mean across images relative to IID. Similar to text-to-image generation, adding SR further improves all cases, indicating that maintaining on-manifold quality increases sample efficiency. Qualitative results are shown in Figure˜4.
@@ 5 Conclusion
We introduced an importance-weighted non-IID sampling framework for flow matching that (i) applies score-based regularization for diversity velocity to push concurrent trajectories apart while keeping them on-manifold, and (ii) computes importance weights along trajectories by learning a residual flow. On a Gaussian mixture, score-regularization improves sample quality while preserves diversity, and our importance weighting yields more accurate expectation estimates than IID or baselines. On Stable Diffusion 3.5 and FLUX.1-Fill, diversity with score-regularization consistently reduces sample representation error.
Limitations and future work.
Although the residual flow $r_{\phi}$ can use a smaller architecture than the base flow, training it at scale remains computationally demanding. Given our current resources, we leave large-scale training of $r_{\phi}$ to future work.
[…]
Supplementary Material
@@ Appendix A Importance Sampling
$\mathbb{E}_{X\sim p}[f(X)]\approx\frac{1}{n}\sum_{i=1}^{n}f\!\left(X^{(i)}\right),\qquad X^{(i)}\overset{\text{IID}}{\sim}p,$ (S1)
@@
$(X^{(1)},\ldots,X^{(n)})\sim p_{\text{joint}}.$ (S2)
@@
$\hat{\mu}=\frac{1}{n}\sum_{i=1}^{n}w\!\left(X^{(i)}\right)\,f\!\left(X^{(i)}\right).$ (S3)
@@
$\mathbb{E}_{(X^{(1)},\ldots,X^{(n)})\sim p_{\text{joint}}}\hat{\mu}=\mathbb{E}_{X\sim p}[f(X)].$ (S4)
@@
$p^{\prime}_{i}(x)=\int p_{\text{joint}}\big(x^{(1)},\ldots,x^{(n)}\big)\Big|_{x^{(i)}=x}\,\mathrm{d}x^{(-i)},$ (S6)
@@
$p^{\prime}(x):=p^{\prime}_{1}(x)=\cdots=p^{\prime}_{n}(x).$ (S7)
@@
$\hat{\mu}=\frac{1}{n}\sum_{i=1}^{n}w\!\left(X^{(i)}\right)\,f\!\left(X^{(i)}\right),\qquad\,w(x)=\frac{p(x)}{p^{\prime}(x)}.$ (S8)
@@
$\mathbb{E}_{(X^{(1)},\ldots,X^{(n)})\sim p_{\text{joint}}}[\hat{\mu}$] (S9)
@@ Appendix B Relationship Between the Score and the Velocity in Rectified Flow
$p_{t}(x)=\mathbb{E}_{X_{1}}\!\left[(1-t)^{-d}p_{0}\!\Big(\tfrac{x-tX_{1}}{1-t}\Big)\right].$ (S12)
@@
Differentiating Eq. S12 and using $\nabla\log p_{0}(z)=-z$ for $\mathcal{N}(0,I)$ gives
$\quad\nabla_{x}\log p_{t}(x)$ (S13)
@@
$m_{1}(x,t):=\mathbb{E}[X_{1}\mid X_{t}=x].$ (S17)
Using Eq. S14 in Eq. S15 gives
$t\,m_{1}(x,t)=x+(1-t)^{2}s(x,t).$ (S18)
@@
where the last equality uses Eq. S14. Multiplying Eq. S19 by $t$ and substituting Eq. S18 gives
$t\,v_{\theta}(x,t)=x+(1-t)s(x,t),$ (S20)
@@ C.1 Evolution at a Fixed Position
$\partial_{t}\tilde{p}_{t}(x)=-\nabla_{x}\cdot\big(\tilde{v}(x,t)\tilde{p}_{t}(x)\big),$ (S23)
@@
$\quad\partial_{t}\log\tilde{p}_{t}(x)$ (S24)
@@
$\partial_{t}\log p_{t}(x)=-\nabla_{x}\cdot v(x,t)-v(x,t)\cdot s(x,t),$ (S25)
@@
$\partial_{t}\log p^{\prime\prime}_{\phi,t}(x)$ $=-\nabla_{x}\cdot\big(v(x,t)+r_{\phi}(x,t)\big)$ (S26)
@@
where $p,s$ denote the original flow and $p^{\prime\prime}_{\phi},s^{\prime\prime}_{\phi}$ denote the marginal under diversity. Subtracting Eq. S26 from Eq. S25 gives
$\partial_{t}\log w_{t}(x)$ $=\partial_{t}\log p_{t}(x)-\partial_{t}\log p^{\prime\prime}_{\phi,t}(x)$ (S27)
@@
which matches Theorem˜1.

### v2

@@ 3.3 Discussions
Since $r_{\phi}$ is trained, one can sample independently from $v+r_{\phi}$ (the ODE in Definition 3), which matches the marginal distribution of diversity-enhanced samples. However, IID batches drawn from $v+r_{\phi}$ are typically less diverse than those generated by the coupled ODEs in Definition 2, which explicitly enhance diversity across concurrent trajectories (see Table 1).
@@
In Theorem 1, we estimate the importance weight at the fixed position for each sample as
@@
which, given $v$ and $v+r_{\phi}$ are rectified flows, further simplifies (Proof in Appendix C.2):
@@
[…]
4 Experiments
We first evaluate on a Gaussian mixture with known density, enabling precise diagnosis of diversity, quality, and estimation of importance weights and expectations. We then assess plug-in benefits on large flow models—Stable Diffusion 3.5 Medium [Esser et al., 2024] for text-to-image generation and FLUX.1-Fill-dev [Labs et al., 2025] for image inpainting—to study distribution coverage in complex image settings. Diversity velocities and implementation details for each experiment are in Appendix D and Appendix E, respectively.
4.1 Experiments on Gaussian Mixture
4.1.1 Sample Diversity and Quality
Setup.
Results.
In Table 1, across all diversity objectives, adding SR (soft or hard) significantly improves quality ($\log p$ increases; RMSE decreases) while preserving joint mode coverage. For example, with DPP, $\log p$ and RMSE improve from $7.16$ and $0.527$ to $19.27$ and $0.253$ under SR-hard, while joint coverage remains almost unchanged ($8.55{\to}8.57$). By contrast, DiverseFlow (DPP+DF) improves quality but reduces diversity ($8.55{\to}6.89$), evidencing a diversity–quality trade-off that our score-based regularization alleviates. Finally, IID sampling from the diversity-enhanced marginal fails to recover joint-diversity benefits (e.g., DPP joint coverage $8.55$ vs. marginal $6.51$, close to original IID $6.51$), justifying our design choice of joint sampling via coupled ODEs (Definition 2) rather than IID from the diversity-enhanced marginal (Definition 3).
@@ Results.
Table 1: Sampling diversity and quality results for the Gaussian mixture experiment. Mode coverage, $\log p$, and RMSE are reported as mean(uncertainty) with 95% confidence intervals. The best-performing method for each metric is shown in bold.
@@
0.073(1) 0.082(2) 0.512(3) 0.526(3) 0.458(3) 0.087(1) 0.077(1)
Table 3: Jensen–Shannon divergence results for estimation. Values are mean(uncertainty) with 95% confidence intervals. indicates the variation of our method that estimates the importance-weight evolution at a fixed, final sample position. “Equal” means using non-IID samples but with equal weights, while “IID” means using additional IID samples. Lower is better; the best method is in bold.
SR $T_{1}$ $T_{2}$ $T_{3}$ $T_{4}$ $T_{5}$ $T_{6}$ Mean
@@
Table 4: Coverage radius results on text-to-image generation. Values are mean(uncertainty) with 95% confidence intervals.
@@ 4.2 Importance Weight & Expectation Estimation
For accurate ground truth, we use a 2D Gaussian mixture where each component $k$ has a weight proportional to $2^{k}$. Non-IID joint sets are sampled using a Harmonic DPP with soft SR. The ground-truth importance weights are computed as the ratio between the original density and the non-IID marginal estimated via Local-Likelihood Density Estimation (LLDE) [Loader, 1996, Hjort and Jones, 1996] (see Appendix F). We train the residual velocity $r_{\phi}$ on 1,0001{,}000 joint sets (each containing $10$ trajectories) and evaluate on 10,00010{,}000 sets. We compare three types of estimators: (i) trajectory-based evolution (ours, Theorem 2); (ii) fixed-position evolution at a terminal state (Theorem 1); and (iii) density baselines—k-Nearest Neighbors Density Estimation (kNN) [Loftsgaarden and Quesenberry, 1965], Gaussian Kernel Density Estimation (KDE) [Silverman, 1986], and Multivariate Gaussian Fit (MGF) [Bishop and Nasrabadi, 2007] (see Appendix G).
@@
We report squared error (SE) and ranking metrics—Kendall’s $\tau_{b}$ [Kendall, 1945], Spearman’s $\rho$ [Spearman, 1987], and graded AP (gAP) [Moffat et al., 2022] (see Appendix H)–computed on the top $50\%$ most-confident ground truth. In Table 2, density baselines exhibit large SE and weak ranking, while the trajectory-based estimator consistently outperforms the fixed-position variant, since path integration avoids out-of-distribution inputs to $r_{\phi}$.
@@
We estimate $\mu{=}\mathbb{E}f(X)$, where $f(x)$ maps each sample to a one-hot vector indicating the index of its nearest Gaussian. Ground-truth $\mu$ is computed from 1,000,0001{,}000{,}000 IID samples drawn from the original flow-matching model, yielding the probability that a generated sample belongs to each Gaussian. We evaluate on joint samples via Jensen–Shannon (JS) divergence [Lin, 1991], which remains well-defined when some predicted probabilities are zero. In Table 3, non-IID samples without weighting are biased and perform worse than IID; fixed-position weighting reduces bias but still trails IID; the trajectory-based estimator achieves the strongest performance.
@@
Table 5: Coverage radius results on image inpainting. Values are mean(uncertainty) with 95% confidence intervals.
@@ 4.3 Experiments on Text-to-Image Generation
We next test Stable Diffusion 3.5 Medium [Esser et al., 2024] for text-to-image generation. We study whether using diverse samples can better cover the IID samples $S_{\text{IID}}$ with a fixed budget of samples $S_{\text{Rep}}$, measured by the coverage radius
@@
Table 4 shows that introducing diversity terms reduces the coverage radius for all prompts $T_{1}$ to $T_{6}$ relative to IID (taking $S_{\text{Rep}}$ as 10 IID samples), indicating improved coverage under a fixed budget. Adding our score-based regularization (SR) further improves performance across all prompts, suggesting that maintaining on-manifold quality increases sample efficiency (fewer “wasted” samples). Qualitative results are shown in Figure 3.
Figure 4: Qualitative results for image inpainting. Two samples are shown from the joint samples, with all methods sharing identical initialization. In Sample 1, DPP introduces artifacts (highlighted by the black rectangle) that are removed by SR (hard). While enhancing quality, DPP+SR retains the diversity of DPP, as illustrated in Sample 2.
@@ 4.4 Experiments on Image Inpainting
We additionally evaluate performance on image inpainting, which is more constrained than text-to-image, to test whether improvements persist. We use FLUX.1-Fill-dev [Labs et al., 2025]. We select 10 square, high-quality images from the Open Images Dataset V7 [Kuznetsova et al., 2020] and apply a mask with four squares to inpaint. For each image, we generate $|S_{\text{IID}}|\!=\!1000$ inpaintings and evaluate $|S_{\text{Rep}}|\!=\!10$ jointly sampled inpaintings (100 repetitions). The coverage radius (Eq. (26)) is computed over average latent features within the masked regions.
Table 5 shows that diversity terms reduce coverage radius for most inputs and improve the mean across images relative to IID. Similar to text-to-image generation, adding SR further improves all cases, indicating that maintaining on-manifold quality increases sample efficiency. Qualitative results are shown in Figure 4.
@@ 5 Conclusion
We introduced an importance-weighted non-IID sampling framework for flow matching that (i) applies score-based regularization for diversity velocity to push concurrent trajectories apart while keeping them on-manifold, and (ii) computes importance weights along trajectories by learning a residual flow. On a Gaussian mixture, score-regularization improves sample quality while preserving diversity, and our importance weighting yields more accurate expectation estimates than IID or baselines. On Stable Diffusion 3.5 and FLUX.1-Fill, diversity with score-regularization consistently reduces sample coverage radius.
@@ Appendix A Importance Sampling
$\mathbb{E}_{X\sim p}[f(X)]\approx\frac{1}{n}\sum_{i=1}^{n}f\!\left(X^{(i)}\right),\quad X^{(i)}\overset{\text{IID}}{\sim}p,$ (27)
@@
$(X^{(1)},\ldots,X^{(n)})\sim p_{\text{joint}}.$ (28)
@@
$\hat{\mu}=\frac{1}{n}\sum_{i=1}^{n}w\!\left(X^{(i)}\right)\,f\!\left(X^{(i)}\right).$ (29)
@@
$\mathbb{E}_{(X^{(1)},\ldots,X^{(n)})\sim p_{\text{joint}}}\hat{\mu}=\mathbb{E}_{X\sim p}[f(X)].$ (30)
@@
$p^{\prime}_{i}(x)=\int p_{\text{joint}}\big(x^{(1)},\ldots,x^{(n)}\big)\Big|_{x^{(i)}=x}\,\mathrm{d}x^{(-i)},$ (32)
@@
$p^{\prime}(x):=p^{\prime}_{1}(x)=\cdots=p^{\prime}_{n}(x).$ (33)
@@
$\hat{\mu}=\frac{1}{n}\sum_{i=1}^{n}w\!\left(X^{(i)}\right)\,f\!\left(X^{(i)}\right),\quad\,w(x)=\frac{p(x)}{p^{\prime}(x)}.$ (34)
@@
$\mathbb{E}_{(X^{(1)},\ldots,X^{(n)})\sim p_{\text{joint}}}[\hat{\mu}$] (35)
@@ Appendix B Relationship Between the Score and the Velocity in Rectified Flow
$p_{t}(x)=\mathbb{E}_{X_{1}}\!\left[(1-t)^{-d}p_{0}\!\Big(\tfrac{x-tX_{1}}{1-t}\Big)\right].$ (38)
@@
Differentiating (38) and using $\nabla\log p_{0}(z)=-z$ for $\mathcal{N}(0,I)$ gives
$\quad\nabla_{x}\log p_{t}(x)$ (39)
@@
$m_{1}(x,t):=\mathbb{E}[X_{1}\mid X_{t}=x].$ (43)
Using (40) in (41) gives
$t\,m_{1}(x,t)=x+(1-t)^{2}s(x,t).$ (44)
@@
where the last equality uses (40). Multiplying (45) by $t$ and substituting (44) gives
$t\,v_{\theta}(x,t)=x+(1-t)s(x,t),$ (46)
@@ C.1 Evolution at a Fixed Position
$\partial_{t}\tilde{p}_{t}(x)=-\nabla_{x}\cdot\big(\tilde{v}(x,t)\tilde{p}_{t}(x)\big),$ (49)
@@
$\quad\partial_{t}\log\tilde{p}_{t}(x)$ (50)
@@
$\partial_{t}\log p_{t}(x)=-\nabla_{x}\cdot v(x,t)-v(x,t)\cdot s(x,t),$ (51)
@@
$\partial_{t}\log p^{\prime\prime}_{\phi,t}(x)$ $=-\nabla_{x}\cdot\big(v(x,t)+r_{\phi}(x,t)\big)$ (52)
@@
where $p,s$ denote the original flow and $p^{\prime\prime}_{\phi},s^{\prime\prime}_{\phi}$ denote the marginal under diversity. Subtracting (52) from (51) gives
$\partial_{t}\log w_{t}(x)$ $=\partial_{t}\log p_{t}(x)-\partial_{t}\log p^{\prime\prime}_{\phi,t}(x)$ (53)
@@
which matches Theorem 1.